### v2

# Experiences with Sub-Arctic Sensor Network Deployment

Abstract
This paper discusses the experiences gained from designing, deploying and maintaining low-power wireless sensor networks in three geothermally active remote locations in Iceland. The network was deployed to assist researchers in collecting soil temperature data which would help them investigate the impact of global warming on (sub)Arctic climate and subsequent carbon release. Functional networks with more than sensor nodes from three sites with no direct access to power and the Internet have been providing researchers insight into the warming impacts since 2021. The network employs low-power primary cell-powered wireless sensor nodes equipped with DASH7 communication protocol and solar-powered DASH7-cellular gateways, providing real-time data and remote access to sensors and devices deployed in the field. We present a detailed discussion of different network components, their architecture, and the network’s overall performance and reliability.
Index Terms:
Wireless Sensor Networks, Environmental Sensing, DASH7, LPWANI Introduction
The Arctic ecosystem is a large reservoir of carbon, developed over millions of years. Climate changes and global warming cause the ecosystem to release a considerable amount of stored carbon, significantly impacting the carbon cycle. It is estimated that close to % of the global carbon is stored in northern high latitude regions, despite it covering only % of global soil surface [1]. Therefore, understanding the future release of carbon from the Arctic system and its impact on climate is crucial. However, estimating the effect of future warming requires soil temperature manipulating experiments that are often costly and not scalable. In this regard, the ForHot site located in the village of Hverageri in southern Iceland [2] offers a geothermally warmed soil. This organic warming in temperature due to geothermal activity can be used as a large-scale controlled environment to study the impact of global warming and future carbon release. Exploiting this natural laboratory and utilizing the technological advancements in Internet of Things (IoT) and Machine Learning (ML), an interdisciplinary research program, FutureArctic was established, in an attempt to investigate the response of the sub(Arctic) ecosystem to climate changes and the release of carbon. To support this investigation, we designed and deployed a Wireless Sensor Network (WSN) in ForHot, which provides real-time access to soil temperature and other physical parameters from three remote locations. The network has been operational since April 2021, with an uptime of over three years to date. This paper outlines the experiences gained from these network deployments and maintenance and discusses the long-term reliability and performance of the network.
II Deployment Sites
The ForHot site is divided into different sections based on the nature of warming as Forest New (FN), Grassland New (GN) and Grassland Old (GO) (Fig. 2) [3]. The soil at the GO site has been warmed since 1963. Whereas, the warming at FN and GN started only after a massive earthquake in 2008 which affected the nearby geothermal systems and consequently started the warming process [3]. Only the GN and GO sites are considered for the deployment and FN is excluded. The GN site is further divided into two as GN 1-3 and GN 4-5. These sites are independent but are collectively named GN as are recently warmed. The GN 4-5 and GO sites are characterised by the presence of thick and tall grass whereas there are spruce trees present at the entrance of the GN 1-3. Additionally, all the sites are characterised by tough terrains and steep hills spread over a wide area. The sites receive heavy snowfall during the winter and can be under snow cover for most of the winter months. The air distance between GN 1-3 and GO is around but there are no connecting paths between the two. The GN 4-5 falls around in the middle of both the sites. It may be noted that while the Google map shows the sites close to each other, they are obstructed by hills.
Each site is divided into (GN/GO 1-5) plots and each plot is further divided into transects (A-F) of x size. These transects are classified based on the warming level of the soil; transect is un-warmed. Taking transect as reference, is warmed approximately by , by , by and by . . It is worth mentioning that these warming levels are based on the measurement campaigns conducted between 2013 and 2016 and the soil temperatures were sampled at a depth of [3].
We were required to measure soil temperature at from total locations (3x6 in GN 1-3, 2x6 in GN 4-5 and 4x6 in GO 1-4), weather condition at GN 4-5 and water content at each site (3).
III Network Components and Design
This section gives an overview of the architecture of the overall network and the hardware-software design of different network components.
III-A Choices of Wireless Communication
We considered three possible Low Power Wide Area Network (LPWAN) standards at the initial stage of design- LoRaWAN, NB-IoT and DASH7. LoRaWAN can provide long-range connectivity at low data rates with extremely low power consumption. NB-IoT, on the other hand, co-exists with cellular networks and can enable a wide range of connectivity. However, NB-IoT has a relatively higher power consumption than LoRaWAN devices [4]. DASH7 is another communication protocol for low power IoT in the ISM band. DASH7 can enable connectivity over in outdoor conditions. Therefore, DASH is a mid-range protocol, not truly LPWAN technology. DASH7 has comparable power consumption levels to LoRa and can provide low-power downlink through an intermittent wake-up protocol [5].
When the FutureArctic deployment was planned, no NB-IoT network was established in Iceland. Consequently, NB-IoT was excluded from further discussions. Out of the remaining two, LoRaWAN has an advantage of long-range over DASH7. Whereas, DASH7 has the advantage of low-power wake-ups which is important for establishing downlink communication to the devices. Though LoRaWAN supports downlink transmission, it is either bound to up-link transmissions (Class A) or requires strict synchronization (Class B). DASH7, on the other hand, requires no periodic synchronization to enable downlink transmission. Moreover, due to the extreme terrain, it was unsure if LoRaWAN would connect all the networks with a single gateway. Additionally, during the initial planning, it was decided that the network would provide real-time ground truth for drone-based remote sensing. Therefore, the low-power bidirectional capability of DASH7 Alliance Protocol (D7A) was considered an advantage. Subsequently, we decided to work with D7A.
III-B Wireless Sensor Nodes
The design of the sensor node is of utmost importance to ensure the long-term and reliable operation of the network. Hardware-software architecture with the lowest possible energy consumption must be employed while providing an interface to a soil temperature sensor (DS18B20: 1-Wire), water content sensor (Meter Group TEROS 10: SDI-12), and weather station (Meter Group ATMOS 41: SDI-12). Moreover, it must be reliable, modular, and easily re-configurable in architecture to provide a seamless sensor interface that anyone without much technical know-how can deploy and work with. This is an important aspect of the design considering that the network will be deployed miles away and troubleshooting and maintenance visits are costly. Therefore, the major requirements of the sensor node design are:
low power consumption to enable long-term operation.
support a variety of sensors with different interfaces such as 1-wire, SDI-12; analog, I2C and SPI for future sensors.
work seamlessly with any sensor. This implies the device should not require reprogramming or reconfiguration to support a different sensor, enabling modular architecture.
The hardware architecture of the sensor nodes is based on the low-power OCTA board111https://shorturl.at/WQvJE. OCTA uses a CMWX1ZZABZ-091 radio modem from Murata which can be programmed as a DASH7 transceiver. There are GPIOs and peripherals connections available through a 10-pin connector on the OCTA board which can be used to connect external peripherals or extension boards (Fig. 3a.a). We use an extension board to provide additional sensors and interfaces (Fig. 3a.b). The extension board houses a temperature-humidity sensor, a watchdog timer, flash memory for data buffering, a configurable power supply ( = , ), an ultra-low power ADC, and a GPIO pins. The configurable power supply allows the device to accommodate external sensors with different voltage requirements. The entire hardware of the node is powered using a primary cell [6].
The sensor nodes run the open-source DASH7 implementation from SubIoT222https://github.com/Sub-IoT to enable DASH7 connectivity along with additional application firmware to support and manage sensors and remote access. The application layer of the node is required to perform three primary tasks (i) sample connected sensors at predefined intervals and transmit the data (ii) provide a seamless interface to sensors and (iii) manage downlink messages and act based on the specific downlink requests. The hardware and software stack of the sensor node is depicted in Fig. 3b. Since D7A is a file-based communication protocol, all the sensors and peripherals on the node can be accessed over-the-air using standard file operations [5]. Therefore, at the application layer, we define a file for the sensor data which can be read to get the current measurement, and sensor configuration which can be accessed to update sensor parameters like sensor type, sampling rate, resolution, etc. The sensor type parameter in the configuration file enables changing the sensor interface without reprogramming the node. The application layer has a software driver for each sensor type (I2C, SDI-12, 1-Wire and SPI). Therefore, updating the sensor type causes the application to load the corresponding driver, even on the run.
The current consumption pattern of the sensor node on wake-up to acquire and transmit data is shown in Fig. 4a. The entire process of data acquisition, processing, and transmission requires around of energy. The sensor node is configured to acquire and transmit packets in every minutes, transmitting packets per day. The application requires only one packet every hour, but an oversampling rate is set to deal with packet losses. On the other hand, the intermittent low-power wake-ups sniffing for incoming packets happen and consume around . The spikes at the left and right sides of the highlighted area in Fig. 4b show these wake-ups. Between all these energy-consuming tasks, the device stays in sleep mode which consumes around . It is also worth noting that only the radio stack of the sensor node is activated during the low-power wake-ups and all the peripherals are in sleep mode to keep the energy consumption minimum.
Using the measurements shown in Fig. 4, the total energy consumption of the node per day is around , with the sleep mode, intermittent wake-ups, and the data actions and transmission , consuming around , , and respectively. Therefore, the ideal lifetime of the device should be around years assuming the entire capacity of the battery is available. However, this is an idealistic calculation considering the battery is exposed to extremely low-temperature conditions which can significantly reduce its performance. In Section V, we discuss the actual lifetime achieved for the nodes.
III-C Gateway
A wireless gateway is required to collect the data from all the sensor nodes deployed and route it to a backend where the data can be stored and accessed as required. Therefore, the gateway should have two network interfaces, one communicating with the low-power sensor nodes and the other for routing data over an internet connection to the backend. Consequently, the power consumption of the gateways is significantly higher than the sensor nodes, making the selection of network interface and power supply crucial. Since there was NB-IoT connectivity available, we had to rely on GPRS for internet access. However, GPRS connection leads to a very high power consumption which cannot be supplied with primary cells. Therefore, we set up a solar power station with solar panel and a battery. In addition, to withstand extreme weather conditions where temperature can reach well below , a battery with self-heating was used. The gateway has an average current consumption of at , giving a backup time of more than five months in ideal conditions.
A picture of the deployed gateway is shown in Fig. 1a. The gateway uses an OCTA board as the DASH7 module and a Quectel BG 96 module as the GPRS modem. Similar to the sensor node, the gateway runs the SubIoT stack configured as a DASH7 gateway. However, the gateway does not implement the application layer, instead, the application layer is transferred to the backend to reduce software complexities.
III-D Network Architecture
The deployment sites are scattered over a wide area with each site divided by hilly terrain. Moreover, none of the sites have direct access to power or internet connectivity. Therefore, we decided to deploy an independent star network at each location as shown in Fig. 5. This implies each site needs a gateway and a solar power station. The gateway connects to the backend which runs the database and visualization framework as depicted in Fig. 5. MQTT enables data transfer between the gateway and the backend, Influxdb stores the collected data and Grafana makes the front-end.
As stated earlier, the DASH7 gateway does not parse the application layer packets. Instead, it forwards the raw packets to the backend. The Application Layer Programming (ALP) of the gateway stack is implemented at the backend in Python using the pyd7a package available from SubIoT. pyd7a allows splitting the DASH7 stack between a low-end and a high-end device; for instance, a low-power microcontroller and a remote virtual machine. This brings another advantage that the backend can talk to the remote sensor nodes in a similar way as the gateway would do; i.e., using file names and node IDs. Therefore, a sensor node in the field can be directly accessed from the backend without bothering about the gateway. The communication between the application layer and the network layer happens over MQTT, wrapping DASH7 packets in MQTT.
IV Deployment
The deployment was planned to be carried out in different phases with each phase covering GN 1-3, GN 4-5 and GO respectively. The incremental deployment was also motivated by the hilly and harsh terrain of the sites which were not easy to access and complete with minimal resources in a single attempt. The first deployment, i.e., at the GN 1-3 was planned to carry out in June 2020, covering the first plots and transects. The deployment would therefore involve soil temperature sensors, one water content sensor and a gateway. Unfortunately, all the plans went in vain with the massive outbreak of COVID-19 and subsequent lockdowns and travel restrictions. Eventually, the first phase of deployment at GN 1-3 was carried out in April-May 2021. The distance between the gateway and the last node in this site is around .
After the successful deployment of the network at GN 1-3, we tried to combine both the GN 4-5 and GO in another attempt in November 2021. This phase intended to deploy soil temperature sensors in GN 4-5 and in GO, a water content sensor in each site and a weather station. Nonetheless, the deployment failed due to adverse weather conditions and reduced daylight hours in the winter. Specifically, moving heavy solar panels, batteries and associated infrastructure for the solar power station became difficult, and hence the deployment was eventually abandoned. However, most of the sensor nodes were already deployed on the field and were left untouched. Later, the gateway and solar power station deployment were completed in July/Aug 2022. After the successful third phase, we had total sensor nodes and gateways active in the field. The gateways in both the GN 4-5 and GO cover nodes in a diameter of approximately .
In summer 2023, all the sensor nodes from GO and some from the other two sites were moved to a new experiment site close to GN 4-5 and they communicate with the GN 4-5 gateway.
V Results and Discussions
A decade back WSN and deployment used to be a challenging task. Ensuring long-term operation and establishing low-power connectivity without depleting batteries created troubles for engineers. However, with the advancements in low-power electronics and connectivity, it has become relatively easy to deploy and maintain WSNs. We have so far achieved more than three years of network lifetime with only two minimal maintenance visits.
The daily averages of the soil temperature data collected are plotted in Fig. 6. Please note that we plotted the data until August 2023, since some nodes were relocated for a different experiment. The annual average temperature for each transect in the entire GN is depicted in Fig. 6c. The warming levels of the soil are visible from the plot. However, the warming levels are different for certain plots, particularly at GN 1-4. We are currently validating the data sets and evaluating the measurements against a set of previously deployed temperature sensors on the field. All the data (soil temperature, air temperature, humidity, weather and RSSI) including the measurements from the weather station and GO site are available publicly333https://www.futurearctic.be/data-access.
Reliability: Multiple parameters can be considered to assess the reliability of a network. Autonomous functioning and the downtime of the network can be primary reliability indicators. As of May 2024, the network at GN 1-3 has reached a total of three years in operation. Similarly, by August 2024, the GN 4-5 has reached two years of operation. We do not consider GO site since it was dismantled after a short duration. Over the three years of operation at GN 1-3, we experienced minimal network downtime that required a field visit. However, there had been multiple short-duration outages lasting a couple of days to weeks mostly due to the dis-functioning of the gateways due to power or connectivity issues. Nonetheless, the gateway was able to recover itself, thus keeping the overall downtime short. Whereas at the GN 4-5, we had two major outages which lasted one month or more. In April-May 2023, the network went down for a month and we received no data. As a result, a field visit was conducted, during which it was discovered that the cellular antennas of the gateways had been left disconnected. Though we are unsure about how this happened, we presume this is due to the strong winds hitting the flaps of the antenna. In another instant, the network became dysfunctional for more than months starting from November 2023. This time, we had a surprising issue as the battery got damaged. Since it was during the winter, moving heavy batteries to the deployment site was nearly impossible, and we had to wait till the snow subsided. Consequently, the network was rebooted only in April 2024. A similar issue was reported at the GO site during the winter of 2023, which further reduced the actual uptime of the GO network. We encountered the same antenna issues at the GO site in November 2022. Following this, the network was not rebooted; instead, all the nodes were relocated.
While network uptime demonstrates the overall reliability of the network, the Packet Reception Ratio (PRR) gives an idea of the node-level reliability. We estimate the PRR of each sensor node as the percentage of the total received packets per day to the expected number of packers, i.e., packets per day. The monthly averages of the PRR for each node for the entire deployment duration are plotted in Fig. 7. The previously discussed major and minor network outages can be observed in the graph as disconnected points. From Fig. 7b, almost all the nodes in GN 4-5 have a high PRR throughout the time, except the months close to the outages. The high PRR can be attributed to the relatively low node density and the area of deployment. The dips in PRR just before the outage could be the early symptoms of the outage which we missed noticing. In addition to that, the winter months also recorded a relatively low PRR for those nodes far from the gateway.
The nodes in GN 1-3 have a rather inferior PRR. We attribute multiple reasons to this behaviour. Firstly, the gateway and the nodes are separated by a greater distance; more than . Additionally, there is a blockage created by spruce trees between the gateway and the site. This blockage is significant during the growing season, i.e., summer. Consequently, a decrease in PRR can be observed for most of the nodes in the summer months along with the snow cover-induced decay in winter. This site has the highest snow accumulation, and all the nodes will be under a snow blanket in the winter, significantly increasing the transmission loss. Overall PRR for all the nodes are shown in Fig. 7c. It can be observed that a couple of nodes have sub-par PRR. However, since we over sampled the sensors, we still received enough data to estimate the temperature accurately.
Battery lifetime: Our initial calculations based on the power profile and battery capacity estimated close to years of lifetime. However, we observed that in real life, this number deviates greatly due to many factors. We have been monitoring the battery voltage of each node, the monthly averages of which are plotted in Fig. 8. The nodes in GN 4-5 still have stable battery voltage, except for a couple of nodes. Whereas the nodes in GN 1-3 have already started showing notable fluctuations in battery voltage. After November 2023, some of the nodes have stopped sending data. We also replaced old batteries on some devices, while re-organizing the network in the summer of 2023. An increase in the voltage of some of the nodes (eg: GN 2A) is due to the battery replacement. While we are unsure how long the other devices in both GN sites would run without a battery replacement, it is certain that, a year lifetime is nearly impossible. We believe it is mostly the temperature that impacts the lifetime. On inspecting the data sheet of the battery [6], the highest capacity is achieved when discharged at around and . A decrease in discharge current, as well as operating temperature, significantly diminishes the capacity. When discharged at and , the achievable capacity is around . Considering that the average annual temperature in Iceland can be less than , a further degradation in capacity may be expected. As of June 2024, we have a total of sensor nodes active. We plan to maintain the network further and closely monitor the lifetime.
Lessons Learned: The process of developing, implementing, and supporting the WSN has provided us with significant knowledge that we think is worth sharing with the research community. While it has become relatively easier to maintain a WSN compared to earlier times, there are still challenges that need to be addressed. Particularly, supporting sensor nodes with longer lifetimes is still a concern. Though we marginally succeeded in resolving this issue using primary cells, the environmental factor can significantly impact the capacity of the primary cells, as we already saw. Not to mention that even the rechargeable battery of our solar bank failed which led to a major outbreak in the service. Therefore, our experience validates the widely discussed inefficiencies of batteries in powering low-power devices, particularly in harsh and hard-to-reach environments.
We were well aware that energy harvesting could have been a potential solution to the battery lifetime issue. However, we were hesitant to depend on it for several reasons: first, there were no off-the-shelf, proven, and reliable energy harvesting solutions available on the market. Countless energy harvesting architectures exist in the literature, but most are experimental. Though we could have experimented with energy harvesters, we deliberately avoided this considering the immediate requirement for data. The availability of high energy density primary cells in small form factor made the choice easy. Another factor in favour of the battery is the lack of reliable ambient energy sources. While solar energy is the preferred option for outdoor deployments, the harsh Icelandic winter weather renders it largely inefficient. Nonetheless, over the course of the project, we have uncovered a unique source of energy for outdoor deployments, which is the temperature difference between the soil and the air. While this energy is particularly available at ForHot due to the geothermal activity, it is present even in places without geothermal heating [7]. One notable feature that sets this energy source apart from solar is its near 24/7 availability, providing power both day and night. However, its power density is often limited, making a multisource energy harvester that combines both solar and soil-air thermal energy the optimal choice for outdoor deployment in harsh conditions. We have advanced the research in this context and have developed prototypes and proof-of-concept devices that work with soil-air thermal energy [8, 7] and multisource energy harvesting architectures [9].
COVID-19 had a substantial impact on development and deployment. In addition to travel restrictions, the inability to access the lab environment significantly delayed the work. Moreover, the global chip shortage nearly jeopardized our hardware architecture due to the unavailability of CMWX1ZZABZ-071 modules. However, we managed to refurbish and reuse OCTA boards from a different project, highlighting the importance of reusable hardware. Additionally, during the lockdowns, we were able to carry out limited small-scale deployments in a controlled outdoor environment. This helped us to uncover some of the inefficiencies of the initial hardware design. For instance, no external watchdog timer was initially included in the design as we relied on the internal one. However, during the test run, a couple of sensor nodes abruptly ceased to function and required a hard restart by removing the battery. Upon inspection, it was discovered that the firmware had entered a faulty state due to bugs in the DASH7 stack. Therefore, an external watchdog was added to later versions to handle any unexpected firmware failures.
VI Conclusion
We have discussed the design, deployment and maintenance of low-power wireless sensor networks for monitoring and sensing at the ForHot sites in Iceland. The network has attained more than three years of lifetime and is still functional. While the primary cells provide a quick solution, it requires further scrutiny to evaluate their long-term performance.